OpenStudy (anonymous):

evaluate x/y for x= 2/3 and y= 5/6

OpenStudy (anonymous):

medals & fans rewarded.

OpenStudy (anonymous):

thanks, but i just needed an example of how to do it. c:

sam (.sam.):

Hmm, \[\frac{x}{y}=\frac{\frac{2}{3}}{\frac{5}{6}}= \frac{2}{3} \div \frac{5}{6}\] Then you flip the sign from dividing to multiplying this chages 5/6 to 6/5 \[\frac{2}{3} \times \frac{6}{5}\] Multiply \[\frac{2\times 6}{3 \times 5}=\frac{12}{15}\] Divide the whole thing by 3 to simplify, you get \[\frac{4}{5}\]